OpenStudy (anonymous):

summation 1 to infinity -> n!/(n!+1) convergent or divergent

OpenStudy (anonymous):

\[\sum_{1}^{\infty} n!/(n!+1)\]

OpenStudy (anonymous):

not sure how to break down the factorials

OpenStudy (anonymous):

all it is asking is to use the limit approaching infinity of a_n and see if divergent or needs further testing (which means it approaches 0)

OpenStudy (anonymous):

well \[\large \lim_{n \to \infty} \frac{n!}{n!+1}=1 \]

OpenStudy (anonymous):

how did you do the factorials?

OpenStudy (anonymous):

can you just divide top and bottom by 1/n!

OpenStudy (anonymous):

\[\large \lim_{n \to \infty} \frac{n!}{n!+1}= \lim_{n \to \infty} \frac{n!}{n!(1+ \frac{1}{n!})}= \lim_{n \to \infty} \frac{1}{1+ \frac{1}{n!}} \]

OpenStudy (anonymous):

\(a_n\) does not converge to zero, thus the series can't converge, the series is divergent.

OpenStudy (anonymous):

ahhh ok that makes sense

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

You're very welcome.